# SHIRE: Enhancing Sample Efficiency using Human Intuition in REinforcement Learning

### v1

@@ Abstract
The ability of neural networks to perform robotic perception and control tasks such as depth and optical flow estimation, simultaneous localization and mapping (SLAM), and automatic control has led to their widespread adoption in recent years. Deep Reinforcement Learning has been used extensively in these settings, as it does not have the unsustainable training costs associated with supervised learning. However, DeepRL suffers from poor sample efficiency, i.e., it requires a large number of environmental interactions to converge to an acceptable solution. Modern RL algorithms such as Deep Q Learning and Soft Actor-Critic attempt to remedy this shortcoming but can not provide the explainability required in applications such as autonomous robotics. Humans intuitively understand the long-time-horizon sequential tasks common in robotics. Properly using such intuition can make RL policies more explainable while enhancing their sample efficiency. In this work, we propose SHIRE, a novel framework for encoding human intuition using Probabilistic Graphical Models (PGMs) and using it in the Deep RL training pipeline to enhance sample efficiency. Our framework achieves $25-78\%$ sample efficiency gains across the environments we evaluate at negligible overhead cost. Additionally, by teaching RL agents the encoded elementary behavior, SHIRE enhances policy explainability. A real-world demonstration further highlights the efficacy of policies trained using our framework.
@@ I Introduction
Among the prevalent machine learning techniques, supervised learning has been the most widely adopted, primarily due to its simplicity of implementation. While supervised learning is simple to implement, its simplicity belies its true cost in terms of data and compute requirements. The training cost of modern supervised learning models is of the order of hundreds of ZFLOPs ($1$ $Zettaflop=10^{21}FLOPs$) [13], with some state-of-the-art models requiring 21 YFLOPs ($1$ $Yottaflop=10^{24}FLOPs$) to train [13]. Supervised training requires large, well-labeled, and well-balanced datasets which are notoriously difficult to generate [14], especially for robotic control tasks. In addition, training time, and cost are directly proportional to dataset size. Thus, larger datasets lead to even larger training overheads, making supervised training infeasible for autonomous robotics applications.
@@
Reinforcement Learning (RL) is a machine learning technique that trains an agent to perform a task through repeated interaction with the environment along with a system of rewards and punishments. A major advantage of RL is that while training requires a large number of interactions with the environment, these are still computationally cheaper than training on large supervised datasets. This makes RL ideally suited for tasks such as autonomous robotics, where datasets are too small due to data generation difficulties or do not exist at all. Recently, deep reinforcement learning, where a neural network is used as the agent, has been successfully applied to autonomous robotics tasks such as dextrous in-hand manipulation [16] and quadrupedal walking [17]. Despite these successes, Deep RL has few drawbacks. Convergence to a good policy is dependent on factors such as random initialization, the quality of the reward function, and the representation granularity of the environment itself. Another shortcoming is the large amount of interactions required to converge to a good policy. Such large data requirement may result in poor sample efficiency in Deep RL algorithms. These shortcomings have traditionally been remedied by supplying the agent with a model of its environment (also called its world). This approach is called Model-Based Reinforcement Learning (MBRL). While MBRL improves sample efficiency significantly, it is subject to instability issues caused by biases in the world model. Previous works such as [18, 19] have shown that exploiting prior knowledge of an environment’s transition dynamics can lead to provably better sample efficiency in Deep RL settings [20, 21, 22]. However, such environmental information may not always be available in the real world.
@@
  • We developed a framework (SHIRE) for encoding human intuition as a Probabilistic Graphical Model (PGM) and use it to enhance the sample efficiency and explainability of Deep RL.
  • Provide experimental evidence to show that our framework improves performance by $25\%$ for simple environments and more than $100\%$ for complex environments.
@@ III-B Policy Optimization Algorithm
where $loss_{policy}$ is the policy loss computed using the clipped advantage function, $loss_{entropy}$ is the degree of uncertainty in the policy, and $loss_{value}$ is the mean squared error between the expected and actual reward. The SHIRE framework adds an additional loss term to this loss term, which is described in the next section.
@@ III-C Intuition Encoding
     $e_{i}\leftarrow PGM(s)$triangleright compute “intuitive” actions using probabilistic inference
     $m_{i}\leftarrow M(e,a)$triangleright compute mismatch vector
@@ IV Experiments
To assess the efficacy of the SHIRE framework, we conduct policy optimization experiments on various standard gym environments. Across environments, we tuned a two-layer vanilla agent with 64 neurons per layer and stored the best result. We then enabled our framework and re-ran the best vanilla experiment on each environment with the same seed for comparison. Thus, the vanilla and SHIRE experiments differ only in the intuition loss scaling coefficient. Table I shows the number of samples (i.e., interactions with the environment) required to attain a solved state both with and without SHIRE. In this section, we describe the intuition encoded in each Intuition Net and provide quantitative analyses of the improvements achieved by using our framework and of the overheads imposed by this framework. Note that we consider an environment to be “solved” when a policy achieves an average reward greater than or equal to the “Solved State Reward” (SSR) for that environment over 100 evaluation episodes.
@@ IV-A4 Swimmer Environment
$intuiion\hskip 2.84526ptloss=f(\theta_{1},\theta_{2})$ (8)
@@ IV-C SHIRE Overheads
The latency overheads incurred by the Intuition Loss computation pipeline are approximately proportional to the size of the Intuition Net. Table II shows the intuition loss computation overhead for all the tested environments. It is immediately apparent that the overhead imposed by SHIRE is negligibly small compared to the average time per environment step, which is of the order of tens of milliseconds. Any increase in training time caused by this overhead is compensated for by SHIRE’s sample efficiency gains, as shown by the gains in wall-clock time shown in Table I.
@@ V Conclusion
This work enables rapid prototyping of RL policies which facilitates efficacy checking for various policy architectures, thereby significantly reducing the RL policy development life cycle. We hope this work spurs further research in efficient and explainable RL, thus enabling the development of robust RL policies aimed at safety-critical operations such as autonomous flight.
Acknowledgement: This work was supported by CoCoSys, one of seven centers in JUMP 2.0, an SRC program sponsored by DARPA

### v3

@@ Abstract
The ability of neural networks to perform robotic perception and control tasks such as depth and optical flow estimation, simultaneous localization and mapping (SLAM), and automatic control has led to their widespread adoption in recent years. Deep Reinforcement Learning (DeepRL) has been used extensively in these settings, as it does not have the unsustainable training costs associated with supervised learning. However, DeepRL suffers from poor sample efficiency, i.e., it requires a large number of environmental interactions to converge to an acceptable solution. Modern RL algorithms such as Deep Q Learning and Soft Actor-Critic attempt to remedy this shortcoming but can not provide the explainability required in applications such as autonomous robotics. Humans intuitively understand the long-time-horizon sequential tasks common in robotics. Properly using such intuition can make RL policies more explainable while enhancing their sample efficiency. In this work, we propose SHIRE, a novel framework for encoding human intuition using Probabilistic Graphical Models (PGMs) and using it in the Deep RL training pipeline to enhance sample efficiency. Our framework achieves $25-78\%$ sample efficiency gains across the environments we evaluate at negligible overhead cost. Additionally, by teaching RL agents the encoded elementary behavior, SHIRE enhances policy explainability. A real-world demonstration further highlights the efficacy of policies trained using our framework.
@@ I Introduction
Among the prevalent machine learning techniques, supervised learning has been the most widely adopted, primarily due to its simplicity of implementation. While supervised learning is simple to implement, its simplicity belies its true cost in terms of data and compute requirements. The training cost of modern supervised learning models is of the order of hundreds of ZFLOPs ($1$ $Zettaflop=10^{21}FLOPs$) [13], with some state-of-the-art models requiring 21 YFLOPs ($1$ $Yottaflop=10^{24}FLOPs$) to train [13]. Supervised training requires large, well-labeled and well-balanced data sets that are notoriously difficult to generate [14], especially for robotic control tasks. In addition, training time and cost are directly proportional to dataset size. Thus, larger datasets lead to even larger training overheads, making supervised training infeasible for autonomous robotics applications.
@@
Reinforcement Learning (RL) is a machine learning technique that trains an agent to perform a task through repeated interaction with the environment along with a system of rewards and punishments. A major advantage of RL is that while training requires a large number of interactions with the environment, these are still computationally cheaper than training on large supervised datasets. This makes RL ideally suited for tasks such as autonomous robotics, where datasets are either too small or do not exist at all due to data generation difficulties. Recently, deep reinforcement learning, where a neural network is used as the agent, has been successfully applied to autonomous robotics tasks such as dextrous in-hand manipulation [16] and quadrupedal walking [17]. Despite these successes, Deep RL has few drawbacks. Convergence to a good policy is dependent on factors such as random initialization, the quality of the reward function, and the representation granularity of the environment. Another shortcoming is the large amount of interactions required to converge to a good policy. Such large data requirement may result in poor sample efficiency in Deep RL algorithms. These shortcomings have traditionally been remedied by supplying the agent with a model of its environment (also called its world). This approach is called Model-Based Reinforcement Learning (MBRL). While MBRL improves sample efficiency significantly, it is subject to instability issues caused by biases in the world model. Previous works such as [18, 19] have shown that exploiting prior knowledge of an environment’s transition dynamics can lead to provably better sample efficiency in Deep RL settings [20, 21, 22]. However, such environmental information may not always be available in the real world.
@@
  • A framework (SHIRE) for encoding human intuition as a Probabilistic Graphical Model (PGM) and using it to enhance the sample efficiency and explainability of Deep RL.
  • Provide experimental evidence to show that our framework improves performance by $25\%$ for simple environments and more than $78\%$ for complex environments.
@@ III-B Policy Optimization Algorithm
where $loss_{policy}$ is the policy loss computed using the clipped advantage function, $loss_{entropy}$ is the degree of uncertainty in the policy, and $loss_{value}$ is the mean squared error between the expected and actual reward. The SHIRE framework adds an additional term to this loss, which is described in the next section.
@@ III-C Intuition Encoding
     $e_{i}\leftarrow PGM(s_{i})$triangleright compute “intuitive” actions using probabilistic inference
     $m_{i}\leftarrow M(e_{i},a_{i})$triangleright compute mismatch vector
@@ IV Experiments
To assess the efficacy of the SHIRE framework, we conduct policy optimization experiments on various standard gym environments. Across environments, we tuned a two-layer vanilla agent with 64 neurons per layer and stored the best result. We then enabled our framework and re-ran the best vanilla experiment on each environment with the same seed for comparison. Thus, the vanilla and SHIRE experiments differ only in the intuition loss scaling coefficient. Table I shows the number of samples (i.e., interactions with the environment) required to attain a solved state both with and without SHIRE. In this section, we describe the intuition encoded in each Intuition Net and provide quantitative analyses of the improvements achieved by using our framework and of the overheads imposed by this framework. Note that we consider an environment to be “solved” when a policy achieves an average reward greater than or equal to the “Solved State Reward” (SSR) for that environment over 100 consecutive evaluation episodes.
@@ IV-A4 Swimmer Environment
$intuition\hskip 2.84526ptloss=f(\theta_{1},\theta_{2})$ (8)
@@ IV-C SHIRE Overheads
The latency overheads incurred by the Intuition Loss computation pipeline are approximately proportional to the size of the Intuition Net. Table II shows the intuition loss computation overhead for all the tested environments. It is immediately apparent that the overhead imposed by SHIRE is negligibly small compared to the average time per environment step, which is typically of the order of tens of milliseconds. Any increase in training time caused by this overhead is compensated for by SHIRE’s sample efficiency gains, as shown by the gains in wall-clock time shown in Table I.
@@ V Conclusion
This work enables rapid prototyping of RL policies, facilitating efficacy checking for various policy architectures, thereby improving the RL policy development life cycle. We hope this work spurs further research in efficient and explainable RL, enabling the development of robust RL policies for safety-critical tasks such as autonomous flight.
Acknowledgement: This work was supported by CoCoSys, one of the seven centers in JUMP 2.0, an SRC program sponsored by DARPA